# 2024_subroc.bib

### v1

@@ B.4 Searching for an Injected Subgroup
For details on the experimental setup, see Section 4.3 in the main text.
@@ B.5 Optimistic Estimates
Table 12: Speedups of median runtimes using optimistic estimates. (Continuation of Table 2 in the main text)
@@
ARL 0    381.7 2.6 2.8
0.1    292.1 1.7 1.5
0.3    283.2 1.7 1.1
1    280.6 1.6 1.2
PR AUC 0    11.8 0.6 2.2
0.1    5.2 0.5 0.6
0.3    5.2 0.5 0.6
1    5.2 0.5 0.6
ROC AUC 0    29.8 0.8 0.8
0.1    3.6 0.6 0.5
0.3    3.7 0.6 0.5
1    3.6 0.7 0.5
Table 13: Median runtimes in seconds for ARL with (“On”) versus without (“Off”) optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$. Subgroups were required to cover at least 20 instances and the search stopped at depth 3. 10 measurements were performed for each configuration.
0 0.1 0.3 1
Dataset Off On Off On Off On Off On
UCI Census-Income (KDD) 6899 723.8 6795 712.9 6877 667.3 6872 592.3
OpenML Adult 177.7 11.6 178.0 13.9 177.9 12.9 178.8 11.5
UCI Bank Marketing 28.8 5.9 28.9 6.9 29.0 6.9 29.2 6.8
UCI Credit Card Clients 523.3 7.5 525.2 10.1 530.3 10.0 531.4 10.0
UCI Mushroom 17.7 0.2 17.8 0.3 17.9 0.3 17.8 0.3
Statlog (German Credit Data) 3.3 1.6 3.3 2.2 3.3 2.2 3.3 2.2
UCI Breast Cancer Wisconsin 0.2 0.1 0.2 0.1 0.2 0.1 0.1 0.1
UCI Credit Approval 0.7 0.5 0.7 0.9 0.7 1.0 0.7 1.0
Table 14: Median runtimes in seconds for ROC AUC with (“On”) versus without (“Off”) optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$. Subgroups were required to cover at least 20 instances and the search stopped at depth 3. 10 measurements were performed for each configuration.
0 0.1 0.3 1
Dataset Off On Off On Off On Off On
UCI Census-Income (KDD) 5469 4785 5499 5391 5609 4814 5477 1429
OpenML Adult 187.2 192.5 187.8 197.7 186.4 182.9 187.6 44.1
UCI Bank Marketing 39.6 21.2 39.8 43.3 39.9 40.6 40.0 26.6
UCI Credit Card Clients 633.1 18.8 647.0 658.6 641.0 633.9 645.6 367.6
UCI Mushroom 20.5 1.0 20.7 7.7 20.6 7.6 20.8 7.7
Statlog (German Credit Data) 5.7 6.1 5.6 7.0 5.6 7.0 5.6 6.9
UCI Breast Cancer Wisconsin 0.2 0.2 0.2 0.3 0.2 0.3 0.2 0.3
UCI Credit Approval 1.0 1.2 1.0 1.6 1.0 1.6 1.0 1.6
Table 15: Median runtimes in seconds for PR AUC with (“On”) versus without (“Off”) optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$. Subgroups were required to cover at least 20 instances and the search stopped at depth 3. 10 measurements were performed for each configuration.
0 0.1 0.3 1
Dataset Off On Off On Off On Off On
UCI Census-Income (KDD) 7007 3645 6923 4149 7014 3502 7147 1522
OpenML Adult 192.4 121.7 194.2 173.6 192.2 114.4 191.8 23.9
UCI Bank Marketing 35.8 28.2 35.9 39.5 35.5 37.1 35.6 23.2
UCI Credit Card Clients 595.4 457.1 599.8 578.2 598.9 405.8 597.4 110.5
UCI Mushroom 24.2 3.3 24.3 7.4 24.3 7.4 24.3 7.2
Statlog (German Credit Data) 4.7 6.0 4.7 7.1 4.8 7.1 4.8 6.7
UCI Breast Cancer Wisconsin 0.2 0.1 0.2 0.3 0.2 0.3 0.2 0.3
UCI Credit Approval 0.9 0.9 0.9 1.6 0.9 1.7 0.9 1.7
Table 16: Speedups of median runtimes reported in Tables 13, 14 and 15.
Performance Measure
Weight ($\alpha=\beta$)
@@
ARL 0 9.5 15.3 4.9 69.4 98.9 2.1 2.5 1.4
0.1 9.5 12.8 4.2 52.2 64.3 1.5 1.6 0.8
0.3 10.3 13.8 4.2 52.8 64.4 1.5 1.1 0.7
1 11.6 15.6 4.3 52.9 64.8 1.5 1.1 0.7
PR AUC 0 1.9 1.6 1.3 1.3 7.3 0.8 1.7 1.0
0.1 1.7 1.1 0.9 1.0 3.3 0.7 0.6 0.6
0.3 2.0 1.7 1.0 1.5 3.3 0.7 0.6 0.6
1 4.7 8.0 1.5 5.4 3.4 0.7 0.5 0.5
ROC AUC 0 1.1 1.0 1.9 33.7 21.6 0.9 0.9 0.9
0.1 1.0 0.9 0.9 1.0 2.7 0.8 0.6 0.6
0.3 1.2 1.0 1.0 1.0 2.7 0.8 0.6 0.6
1 3.8 4.3 1.5 1.8 2.7 0.8 0.6 0.7
Table 17: Share of remaining evaluated subgroups in the same searches reported in Tables 13, 14 and 15
Performance Measure
Weight ($\alpha=\beta$)
UCI Census-Income (KDD)
OpenML Adult
UCI Bank Marketing
UCI Credit Card Clients
UCI Mushroom
Statlog (German Credit Data)
UCI Breast Cancer Wisconsin
UCI Credit Approval
ARL 0 4.51% 2.47% 6.34% 0.41% 0.21% 28.23% 20.76% 48.52%
0.1 4.10% 2.37% 6.34% 0.41% 0.21% 28.23% 20.76% 53.70%
0.3 3.74% 2.13% 6.34% 0.41% 0.21% 28.23% 20.76% 61.64%
1 3.12% 1.81% 6.34% 0.41% 0.21% 28.23% 20.76% 67.35%
PR AUC 0 42.38% 48.05% 45.50% 59.99% 8.48% 100.00% 20.76% 51.82%
0.1 48.67% 76.41% 89.06% 87.05% 16.68% 100.00% 51.18% 88.98%
0.3 39.40% 42.41% 75.48% 47.02% 16.70% 100.00% 51.18% 94.74%
1 11.98% 4.73% 24.25% 7.79% 15.56% 89.35% 62.44% 91.12%
ROC AUC 0 83.00% 97.49% 47.80% 3.10% 2.74% 100.00% 71.65% 94.74%
0.1 91.62% 97.49% 98.05% 99.03% 22.29% 100.00% 71.65% 94.74%
0.3 79.95% 87.01% 82.80% 94.49% 22.29% 100.00% 71.65% 94.74%
1 13.92% 11.33% 28.98% 36.66% 22.29% 94.27% 71.65% 92.11%
@@
Figure 7: Histogram of standard deviations divided by medians for all runtimes of Tables 13, 14 and 15. Most series of measurements did not experience noticeable runtime fluctuations compared to their median runtime.
Table 18: Median runtimes in seconds for ARL with (“On”) versus without (“Off”) optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$. Subgroups were required to cover at least 20 instances and the search stopped at depth 4. 10 measurements were performed for each configuration, except for the dataset “UCI Census-Income (KDD)” for which it were 3 measurements.
0 0.1 0.3 1
Dataset Off On Off On Off On Off On
UCI Census-Income (KDD) 98310 5346 98206 5265 98876 4958 98284 4223
OpenML Adult 978.1 20.6 995.1 25.2 969.7 24.0 983.5 22.0
UCI Bank Marketing 119.5 12.4 119.2 15.5 119.9 15.7 119.4 15.6
UCI Credit Card Clients 4510 8.4 4479 11.6 4449 11.6 4491 11.6
UCI Mushroom 109.5 0.3 110.8 0.4 109.8 0.4 109.9 0.4
Statlog (German Credit Data) 9.8 3.8 10.0 5.9 9.9 5.9 9.8 6.0
UCI Breast Cancer Wisconsin 0.2 0.1 0.2 0.2 0.2 0.2 0.1 0.1
UCI Credit Approval 1.7 1.0 1.8 2.2 1.8 2.5 1.8 2.8
Table 19: Median runtimes in seconds for ROC AUC with (“On”) versus without (“Off”) optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$. Subgroups were required to cover at least 20 instances and the search stopped at depth 4. 10 measurements were performed for each configuration, except for the dataset “UCI Census-Income (KDD)” for which it were 3 measurements.
0 0.1 0.3 1
Dataset Off On Off On Off On Off On
UCI Census-Income (KDD) 73835 731.2 72673 73204 74144 57059 74759 11694
OpenML Adult 961.8 1076 965.9 1138 971.8 1014 976.0 116.3
UCI Bank Marketing 150.1 69.9 150.5 199.2 149.8 170.8 148.0 72.0
UCI Credit Card Clients 4907 58.6 5000 5256 4954 4569 4967 1267
UCI Mushroom 107.2 3.6 107.6 29.7 108.3 29.6 106.9 29.9
Statlog (German Credit Data) 15.4 18.2 15.4 25.0 15.2 25.2 15.3 23.4
UCI Breast Cancer Wisconsin 0.2 0.2 0.2 0.4 0.2 0.4 0.2 0.4
UCI Credit Approval 2.3 3.1 2.3 4.8 2.3 4.8 2.3 4.7
Table 20: Median runtimes in seconds for PR AUC with (“On”) versus without (“Off”) optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$. Subgroups were required to cover at least 20 instances and the search stopped at depth 4. 10 measurements were performed for each configuration, except for the dataset “UCI Census-Income (KDD)” for which it were 3 measurements.
0 0.1 0.3 1
Dataset Off On Off On Off On Off On
UCI Census-Income (KDD) 99395 34225 101272 52252 101602 37638 101372 8896
OpenML Adult 1044 444.9 1060 917.9 1032 483.9 1031 52.0
UCI Bank Marketing 137.4 79.2 137.4 189.0 137.7 162.6 138.8 62.3
UCI Credit Card Clients 4779 1899 4823 4199 4842 1681 4731 209.3
UCI Mushroom 146.4 12.5 149.2 29.0 149.0 28.9 148.7 28.5
Statlog (German Credit Data) 13.2 20.8 13.3 28.1 13.3 27.9 13.4 25.3
UCI Breast Cancer Wisconsin 0.2 0.1 0.2 0.3 0.2 0.3 0.2 0.4
UCI Credit Approval 2.1 1.8 2.1 5.1 2.1 5.4 2.1 5.2
Table 21: Share of remaining evaluated subgroups in the same searches reported in Tables 18, 19 and 20
Performance Measure
Weight ($\alpha=\beta$)
UCI Census-Income (KDD)
OpenML Adult
UCI Bank Marketing
UCI Credit Card Clients
UCI Mushroom
Statlog (German Credit Data)
UCI Breast Cancer Wisconsin
UCI Credit Approval
ARL 0 2.39% 0.68% 2.66% 0.04% 0.03% 19.99% 18.97% 27.64%
0.1 2.17% 0.66% 2.66% 0.04% 0.03% 19.99% 18.97% 32.30%
0.3 1.97% 0.61% 2.66% 0.04% 0.03% 19.99% 18.97% 39.50%
1 1.54% 0.53% 2.66% 0.04% 0.03% 19.99% 18.97% 47.98%
PR AUC 0 25.80% 27.77% 22.54% 22.70% 5.31% 99.60% 18.97% 30.55%
0.1 42.09% 66.38% 84.34% 65.46% 10.43% 99.60% 47.47% 83.66%
0.3 27.81% 28.28% 65.27% 17.37% 10.49% 98.95% 47.47% 92.61%
1 4.32% 1.57% 11.72% 1.23% 9.85% 87.02% 57.76% 88.14%
ROC AUC 0 0.99% 96.32% 38.06% 1.05% 2.10% 99.60% 68.17% 93.91%
0.1 90.86% 96.32% 97.18% 92.96% 16.28% 99.60% 68.17% 93.91%
0.3 67.43% 83.35% 74.42% 75.78% 16.28% 99.60% 68.17% 93.91%
1 8.41% 4.76% 15.70% 11.74% 16.28% 92.08% 68.17% 89.20%
Figure 8: Histogram of standard deviations divided by medians for all runtimes of Tables 18, 19 and 20. Most series of measurements did not experience noticeable runtime fluctuations compared to their median runtime.
For the speedups of median runtimes reported in Tables 18, 19 and 20, see Table 2 in the main text.

### v2

@@ B.4 Searching for an Injected Subgroup
For details on the experimental setup, see Section˜4.3 in the main text.
@@ B.5 Optimistic Estimates
Table 13: Speedups of median runtimes using optimistic estimates. (Continuation of Table˜3 in the main text)
@@
ARL 0    617.8 132.4 5.0
0.1    410.1 85.5 2.8
0.3    405.6 84.2 2.6
1    403.0 85.6 2.7
PR AUC 0    699.8 84.4 4.8
0.1    211.9 62.3 2.9
0.3    89.7 85.9 2.9
1    78.6 85.7 2.9
ROC AUC 0    337.7 87.2 8.3
0.1    268.6 54.8 5.0
0.3    277.1 53.4 4.3
1    249.6 68.9 4.3
Table 14: Median runtimes in seconds for ARL without optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$. Subgroups were required to cover at least 20 instances and the search stopped at depth 4. 10 measurements were performed for each configuration, except for the dataset “UCI Census-Income (KDD)” for which it were 3 measurements.
Dataset 0 0.1 0.3 1
UCI Census-Income (KDD) 114771.7 115310.2 114248.1 114822.9
OpenML Adult 1058.0 1056.0 1044.8 1061.8
UCI Bank Marketing 127.8 128.6 128.7 128.9
UCI Credit Card Clients 4826.9 4803.8 4867.5 4800.9
UCI Mushroom 119.2 119.4 118.5 119.1
Statlog (German Credit Data) 10.6 10.7 10.6 10.8
UCI Breast Cancer Wisconsin 0.1 0.1 0.1 0.1
UCI Credit Approval 1.9 1.8 1.8 1.9
Table 15: Median runtimes in seconds for ROC AUC without optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$. Subgroups were required to cover at least 20 instances and the search stopped at depth 4. 10 measurements were performed for each configuration, except for the dataset “UCI Census-Income (KDD)” for which it were 3 measurements.
Dataset 0 0.1 0.3 1
UCI Census-Income (KDD) 84443.9 85060.0 85223.8 84915.9
OpenML Adult 1055.2 1048.9 1036.1 1051.5
UCI Bank Marketing 160.0 159.6 160.2 161.4
UCI Credit Card Clients 5304.0 5305.0 5293.7 5275.4
UCI Mushroom 118.4 117.7 117.4 118.9
Statlog (German Credit Data) 16.2 16.5 16.4 16.5
UCI Breast Cancer Wisconsin 0.3 0.3 0.3 0.3
UCI Credit Approval 2.5 2.4 2.4 2.4
Table 16: Median runtimes in seconds for PR AUC without optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$. Subgroups were required to cover at least 20 instances and the search stopped at depth 4. 10 measurements were performed for each configuration, except for the dataset “UCI Census-Income (KDD)” for which it were 3 measurements.
Dataset 0 0.1 0.3 1
UCI Census-Income (KDD) 118601.6 118895.8 118370.9 118797.9
OpenML Adult 1123.3 1134.3 1127.3 1124.1
UCI Bank Marketing 145.4 146.4 147.7 146.9
UCI Credit Card Clients 5139.3 5138.0 5117.3 5122.6
UCI Mushroom 161.0 162.8 162.8 162.6
Statlog (German Credit Data) 14.2 14.8 14.2 14.1
UCI Breast Cancer Wisconsin 0.2 0.2 0.2 0.2
UCI Credit Approval 2.2 2.2 2.2 2.2
Table 17: Share of remaining evaluated subgroups in the same searches reported in Tables˜14, 15 and 16
@@
Performance
Measure Weight ($\alpha=\beta$)
ARL 0 6.62% 3.82% 0.02% 4.55% 0.03% 0.13% 6.68% 0.48%
0.1 6.07% 3.48% 0.02% 4.55% 0.03% 0.13% 6.68% 0.48%
0.3 5.20% 2.82% 0.02% 4.55% 0.03% 0.13% 6.68% 0.48%
1 3.21% 1.07% 0.02% 4.55% 0.03% 0.13% 6.68% 0.48%
PR AUC 0 42.47% 47.13% 0.02% 17.65% 0.03% 0.13% 6.68% 0.48%
0.1 42.47% 60.95% 0.02% 42.01% 0.07% 0.13% 6.68% 0.48%
0.3 42.47% 63.18% 0.02% 42.77% 0.20% 0.13% 6.68% 0.48%
1 19.63% 7.85% 0.02% 25.62% 0.25% 0.13% 6.68% 0.48%
ROC AUC 0 61.26% 73.32% 0.02% 42.99% 0.04% 0.13% 7.15% 0.48%
0.1 61.26% 73.32% 0.02% 42.99% 0.04% 0.13% 7.15% 0.48%
0.3 61.29% 73.32% 0.02% 42.99% 0.04% 0.13% 7.15% 0.48%
1 26.83% 9.89% 0.02% 36.11% 0.04% 0.13% 7.15% 0.48%
@@
Figure 8: Histogram of standard deviations divided by medians for all runtimes of Tables˜14, 15 and 16. Most series of measurements did not experience noticeable runtime fluctuations compared to their median runtime.